Find the number of computer units x that needs to be sold to produce a maximum profit p ( in pesos) if p(x) =900x-0.1x^2
Answer by rothauserc(4718)

p(x) = 900x - 0.1x^2
this is a quadratic equation whose graph is a parabola that curves downward,
therefore the vertex of the parabola will give us the number of computer units we need to sell to get maximum profit in pesos
if you have not studied calculus, then x is given by
x = -b/2a where b = 900 and a = -0.1
x = -900 / (2*(-0.1)) = 4500 units
if you have studied calculus, then we take the first derivative and set it equal to 0 and solve for x
0 = 900 - (2*0.1)x
x = -900 / (-2*0.1) = 4500 units